On June 21, 1994 at 8:59 GMT, the HRI high voltage was increased from level 12 to 13 in order to approximately recover the detector's original PHA distribution. Figure 26 shows an observation of the on-board UV calibration lamp taken during the HV increase. It exhibits an increase of 4% in the calibration lamp count rate, indicating the HV level change has approximately recovered the original UV sensitivity. In figure 27 the PHA distributions before and after the increase are compared with an observation taken shortly after launch, showing that the mean PHA has recovered to within 0.5 channels of the original. See also Figures 21 and 22.
The date and time given above for the change of HV from step 12 to 13 should not be used ``blindly'' because step 12 was commanded shortly afterwards by the inadvertent use of a standard parameter file. Lacking precise archival records, we recommend that the user examine the temporal status indicator record for any observation taken in June and early July of 1994 if he needs to know the HV setting. In figure 22, notice the ``overlap'' of coverage during this period, with data at both high voltage steps for N132D.
While turning the HRI back into the focus of the ROSAT XRT on March 3, 1997 (after the PSPC intermezzo), the initial high voltage setting (i.e. that used at the beginning of the mission, step 12) was inadvertently re-initiated. The mistake was only recognized two months later and the HV was reset to step 13 (which has been used since 1994 Jun) on 1997 May 6 at 12:40 UT.
As can be seen from Figure 22, the gain is characterized by a quasi-linear decrease with time. Therefore, it was decided to again increase the gain: from step 13 to 14. Tests were carried out in late November and early December 1997 (calibration observations at both steps 13 and 14) to make sure that the higher voltage did not induce an outbreak of new hotspots etc. Step 14 became the standard setting on 1998 Dec 8.
In the course of the annual time reset the old HV setup (step 13) was commanded by mistake and was effective between 10:45 UT on 1998 Jan 20 until 1998 Jan 21 at 10:40 UT. Thereafter things went normal again.